OpenStudy (anonymous):

Separable ODE: xy'=y^2+y

OpenStudy (anonymous):

It says in the problem to set y/x=u I tried that and got: y'=y^2/x+y/x =uy+u since y/x=u, y=ux and y'=u'x+u equating y', I get: uy+u=u'x+u uy=u'x I'm stuck here, can't integrate since there are 3 variables.

OpenStudy (anonymous):

\[x \frac{ dy }{dx }=y ^{2}+y\] in the statement variable separable \[\frac{ dy }{y \left( y+1 \right) }=\frac{ dx }{x }\] make partial fractions of L.H.S and integrate

OpenStudy (anonymous):

That would be one way of solving it, but the book wants me to use substitution, setting u=y/x

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=xy ^{2}+xy\] \[\frac{ dy }{ dx }-xy=xy ^{2}\] \[divide by y ^{2}\] \[y ^{-2}\frac{ dy }{ dx }+y ^{-1}x=x\] \[substitute y ^{-1}=u,-y ^{-2}\frac{ dy }{dx}=\frac{ du }{dx }\] it serves your purpose?

OpenStudy (anonymous):

\[u=\frac{ y }{x }ory=ux,\frac{ dy }{dx }=u+x \frac{ du }{dx }\] \[x \left( u+x \frac{ du }{dx} \right)=u ^{2}x ^{2}+ux\] \[u+x \frac{ du }{dx }=u ^{2}x+u\] i think now you can solve, i have told you many ways.